OpenStudy (anonymous):

PLEASE Solve the compound inequality 3x + 6 > 36 and 3x + 6 less than or greater to 54

OpenStudy (anonymous):

\[36 <3x+6 \le 54\] Is this what you mean?

OpenStudy (amistre64):

i dunno, i wanna see you go for the less than or greater to option ;)

OpenStudy (anonymous):

I mean: 3x + 6 > 36 and 3x + 6 <= 54

OpenStudy (anonymous):

ha ha ha, yep thats a new one on me :D

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

ok first thing to do is to subtract the 6 from each of the terms so \[30<3x \le 48\] now divide by the 3 n 3x to get \[10 < x \le 16\]

OpenStudy (anonymous):

Thank YOu!

OpenStudy (anonymous):

u look like u have already got ur answer.... lol